# Exoplanets synchronization: Learning from Venus’ retrograde rotation

### v1

@@ 4.1 Smooth transitions in a toy model
Figure 10: : Left Solution of Eqn. (14) along a path defined by $b=\Gamma/64$ crossing the regions 0 and I of Fig. 9 before reaching the boundary $b=1$. Right The same with a logarithmic scale. Also shown: The locus of the stationary root $y_{0}=\nu_{0}/\gamma^{\prime}$ (dashed line), and the derivative $dy/dt$ (black).
@@
The toy model can be improved by taking into account that the atmosphere is formed by the planet’s outgassing. So, mass is being transferred from the inner parts of the planet to the atmosphere. This process increases the moment of inertia of the system. The effect of a small uniform increase in the moment of inertia of the planet is one small subtracting quantity Z-Z in $dy/dt$. Hence,
@@
Fig. 11 shows the initial part of one solution for this equation obtained with the same parameters and initial conditions as before for $Z=10^{-3}$. In this solution, $y$ becomes negative as soon as the solution approaches the synchronous stationary solution, still inside the domain 0. The rotation becomes sub-synchronous (i.e. $\nu<0$ , or $\Omega<n$) and later on becomes retrograde. We stress the fact that what is being observed in Fig. 11 is not the transition from prograde to retrograde, but the transition from super-synchronous ($\Omega>n$) to sub-synchronous ($\Omega>n$). But as $|\nu|$ increases, it can become larger than $2n$ and, hence, $\Omega<0$.
@@ 5 Conclusion
The main conclusion of this report is that no collision with other bodies is necessary to convert the rotation of an Earth or super-Earth with an important atmosphere formed in the course of its evolution to retrograde. It is enough for the planet to be at a distance from the host star short enough to allow the tidal torques to almost synchronize the rotation of the planet before the formation of the bulk of its atmosphere (but not so close that the formed atmosphere is destroyed by the stellar radiation). As the atmosphere is formed from the outgassing of the planet, the transfer of mass from the planet to its atmosphere may be responsible for a loss of angular velocity, so that the rotation is subsynchronous when the atmospheric torques become more important than the tidal torques  [^5] Keep in mind that subsyncronous means $\nu<0$ and retrograde means $\Omega<0$. In the continuation, the prevailing atmospheric torques increase the relative rotation of the planet and create a runaway increase of its value, allowing the rotation to become retrograde later on. None of these processes involves a catastrophic event (even if the transformation of the stable stationary solution into an unstable one is one typical example of René Thom’s catastrophes – see Arnold, 1984). The formation of the planetary atmosphere is a continuous, smooth process, which may be more or less efficient but is not a low-probability event.
@@
The example constructed using a toy model in Sec. 4.1 introduced one evolution history of the planet in which the parameters $b$ and $\Gamma$ grow uniformly. This choice is only justified by its simplicity. The mathematical phenomena shown in that example do not depend on the adopted path. The essential points are that the system must remain in domain 0 long time so that it reaches an almost synchronization and then crosses the boundary from domain 0 to domain I where the nature of the synchronous solution changes from stable to unstable. The change to retrograde is not a necessary fate. It depends on the sign of $\nu$ just before crossing. If the motion is still super-synchronous, it will remain prograde forever. Conversely, if it is subsynchronous, it will evolve towards a retrograde rotation.
@@
The dense atmosphere of the planet plays a role not only in the inversion of the planetary rotation, but also in keeping it retrograde. The study of a naked Venus shows that the current retrograde rotation would become prograde again in the absence of the atmospheric torque in less than one million years. The full 3D creep tide theory indicates that, at this inversion, we do not have a simple reversal of the rotation but a continuous migration of the rotation poles, which are displaced until the inversion is completed. The same simulations show that the equinox precession of Venus is not a circulation like the Earth’s, but rather an oscillation.
@@
Acknowledgements This investigation was sponsored by CNPq (Proc. 303540/2020-6) and FAPESP (Proc. 2016/13750-6 ref. PLATO mission). Part of this research was conducted at the Southwest Research Institute in Boulder.

### v2

@@ 4.1 Smooth transitions in a toy model
Figure 10: : Left Solution of Eqn. (14) along a path defined by $b=\Gamma/64$ crossing the regions 0 and I of Fig. 9 before reaching the boundary $b=1$. Right The same with a logarithmic scale. Also shown are the locus of the stationary root $y_{0}=\nu_{0}/\gamma^{\prime}$ (dashed line), and the derivative $dy/dt$ (black).
@@
The toy model can be improved by taking into account that the atmosphere is formed by the planet’s outgassing. So, mass is being transferred from the planet’s inner parts to the atmosphere. This process increases the system’s moment of inertia. The effect of a small uniform increase in the planet’s moment of inertia is a small subtracting quantity Z-Z in $dy/dt$. Hence,
@@
Fig. 11 shows the initial part of one solution for this equation obtained with the same parameters and initial conditions as before for $Z=10^{-3}$. In this solution, $y$ becomes negative as soon as the solution approaches the synchronous stationary solution, still inside the domain 0. The rotation becomes sub-synchronous (i.e. $\nu<0$ , or $\Omega<n$) and, later on, becomes retrograde. We stress the fact that what is being observed in Fig. 11 is not the transition from prograde to retrograde, but the transition from super-synchronous ($\Omega>n$) to sub-synchronous ($\Omega<n$). But as $|\nu|$ increases, it can become larger than $2n$ and, hence, $\Omega<0$.
@@
Figure 12: Same as Fig. 11 along the path $b=2.5\Gamma$ that drives the solution from domain 0 to domain II, instead of domain I, and eventually to the white domain where the synchronous rotation is unstable.
Fig. 12 shows an alternative path, in which $b$ grows more quickly than $\Gamma$ ($b-2,5\Gamma$). While in the previous case, the increase in the angle $\sigma^{\prime}$ led the evolution, here that angle changes slowly, and the evolution is ruled by the increasing coefficient $A^{\prime}$, which means stronger atmospheric tides. The path goes from domain 0 to domain II (instead of domain I), but later on it crosses the boundary $b=\Gamma^{-2}$ and enters the white region where no stable stationary solution exists. $y$ remains near the synchronous solution for a while, but eventually starts a runaway growth to negative values. The rotation becomes sub-synchronous (i.e. $\nu<0$ , or $\Omega<n$) and evolves towards a retrograde one.
@@ 5 Conclusion
The main result of this report is that no collision with other bodies is necessary to convert the rotation of an Earth or super-Earth with an important atmosphere formed in the course of its evolution into a retrograde rotation. It is enough for the planet to be at a distance from the host star short enough to allow the tidal torques to almost synchronize the rotation of the planet before the formation of the bulk of its atmosphere (but not so close that the stellar radiation destroys the formed atmosphere). As the atmosphere is formed from the outgassing of the planet, the transfer of mass from the planet to its atmosphere may be responsible for a loss of angular velocity, so that the rotation is sub-synchronous when the atmospheric torques become more important than the tidal torques  [^5] Keep in mind that sub-synchronous means $\nu<0$ and retrograde means $\Omega<0$. In the continuation, the prevailing atmospheric torques increase the relative rotation of the planet, allowing the rotation to become retrograde later on. None of these processes involves a catastrophic event (even if the transformation of a stable stationary solution into an unstable one is a typical example of René Thom’s catastrophes – see Arnold, 1984). The formation of the planetary atmosphere is a continuous, smooth process, which may be more or less efficient but is not a low-probability event.
@@
The example constructed using a toy model in Sec. 4.1 introduced one evolution history of the planet in which the parameters $b$ and $\Gamma$ grow uniformly. This choice is only justified by its simplicity. The mathematical phenomena shown in that example do not depend on the adopted path. The essential points are that the system must remain in domain 0 for a long time so that it reaches an almost synchronization and then crosses the boundary from domain 0 to domain I, where the nature of the synchronous solution changes from stable to unstable. The change to retrograde is not a necessary fate. It depends on the sign of $\nu$ just before crossing. If the motion is still super-synchronous, it will remain prograde forever. Conversely, if it is sub-synchronous, it may evolve towards a retrograde rotation.
The conclusions come from simplified models, but they are very robust and should not be changed if different models are used (see Correia et al. 2003a,b). The gravitational tidal torque creates one stationary synchronous solution at $\nu\simeq 0$ as long as the eccentricity of the orbit is kept small. The existence of a contrary torque maintains the current retrograde rotation of Venus. Improvements in the atmospheric torque model should not change the fact that, acting alone, it creates an unstable synchronous solution. When the two torques act together, they lead to a synchronous solution that is stable as long as the gravitational tidal torque dominates and unstable otherwise. The result is a mere consequence of these model-independent facts.
@@
The dense atmosphere of the planet plays a role not only in the inversion of the planet’s rotation, but also in keeping it retrograde. The study of a naked Venus shows that the current retrograde rotation would become prograde again in the absence of the atmospheric torque in less than one million years. The full 3D creep tide theory indicates that, at this inversion, we do not have a simple reversal of the rotation but a continuous migration of the rotation poles, which are displaced until the inversion is completed. The same simulations show that the equinox precession of Venus is not a circulation like the Earth’s, but rather an oscillation.
@@
Acknowledgements This investigation was sponsored by CNPq (Proc. 303540/2020-6) and FAPESP (Proc. 2016/13750-6 ref. PLATO mission). Part of this research was conducted at the Southwest Research Institute in Boulder. I thank R. Alves-Silva for his suggestions and comments on an early version of this report.